Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 332     Entry time: 06 Sep 2023, 06:35
Author: Giuseppe Mercurio 
Group: SCS 
Subject: Giuseppe, Justine 

achievements:

- we found optimum hRIXS settings for 1020 eV but moving the detector arm

- we found optimum hRIXS settings for 920 eV but moving the detector arm, and then grating pitch and detect arm

- energy calbibration for 920 eV

problems:

- lost internect connection around 6 am

- PI-MTE3 camera claims an exposure of 30 sec but in reakity it"d more lik r13 sec.

- if FEL intensity goes to 0 i spikes, the jupyter notebook give errors, referring to objecests ofdifferent sie.

plans:

-

ELOG V3.1.4-7c3fd00